What is the length of a diagonal of a square with a side that measures 9 meters?
Thank you very much!

Found 2 solutions by richwmiller, checkley77:
Answer by richwmiller(17219)

We are talking about a right triangle with both legs 9 meters
2a^2=c^2
2*81=c^2
162=c^2
c is a little more than 12 and not quite 13
12.7279
144=12^2
169=13^2

Answer by checkley77(12844)

9^2+9^2=x^2
81+81=x^2
162=x^2
x=sqrt182
x=13.49 m. ans. for the diagonal.